Students are required to answer the following:

A) A company is manufacturing product X and the selling price is $1,500, while the variable cost for each unit is $500, mean while the fixed cost of the operation will be $50,000. Calculate the break-even in units. 1 mark
B) Define Return on capital employed and why it is used. 2 marks
Answer #1

Answer- (a)

Break-even point (units) = Fixed cost/ (Selling price per unit- variable cost per unit)

Break-even point = 50000/(1500-500) =50000/1000

Break-even point = 50 units.

Answer - (b)

Return on capital employed- It is financial ratio which indicate a company's capital efficiency and profitability. The formula for return on capital employed = Earning before interest and tax/ (Total assets - Total current liabilities). This ratio helps in understanding about how well a company is generating profits from its capital resources.